OpenStudy (anonymous):

Question: 18 of 27: Given the ordered pair (3, y), what value of y would make the ordered pair a solution of the equation 4x − 2y = 24? Select one of the options below as your answer: * A. 12 * B. 6 * C. -6 * D. -8

OpenStudy (anonymous):

we need to plug in 3, and then solve for y: 4*3-2y=24 12-2y=24 -2y=12 y=-6 C) -6

OpenStudy (anonymous):

when someone answers a question for you, it's polite to click on "Good Answer" to remove it from the unanswered questions list and give the appropriate person credit for answering. :)